OpenStudy (anonymous):

[Part 1 (4 points)] Show all work in dividing 9x^2-1/8x-4 / 3x^2+5x-2/2x^2+1x-1 (list restrictions). [Part 2 (2 points)] Use complete sentences to explain your process. (

OpenStudy (campbell_st):

factor all the equations 1st 9x^2 -1 = (3x -1)(3x +1) 8x - 4= 4(2x -1) 3x^2 + 5x - 2 = (3x -1)(x + 2) 2x^2 + x - 1 = (2x -1)(x +1)

OpenStudy (campbell_st):

the rule for dividing by a fraction is to flip and multiply \[\frac{(3x -1)(3x+1)}{4(2x -1)} \times \frac{(2x -1)(x +1)}{(3x-1)(x + 2)}\] cancel the common factors then expand for the solution